OpenStudy (anonymous):

If y=x3+5x and dx/dt=5, find dy/dt when x=2.

OpenStudy (anonymous):

ok for this question i think im doing something wrong i got the answer of 17 but this is not right please help

OpenStudy (nubeer):

can you find dy/dx of the given equation?

OpenStudy (anonymous):

dy/dt=3x^2+5

OpenStudy (phi):

when you take the derivative wrt to t you should get a dx/dt term

hartnn (hartnn):

u need to use chain rule d/dt(5x) = 5 d/dx (dx/dt)

OpenStudy (phi):

instead of d/dx x^3 = 3 x^2 dx/dx you do d/dt x^3 = 3 x^2 dx/dt

OpenStudy (phi):

if you think \[ \frac{d}{dx} x= 1\] think this \[ \frac{ d}{dx} x= \frac{dx}{dx}= 1\]

OpenStudy (anonymous):